# FISH-Tuning: Enhancing PEFT Methods with Fisher Information

### v1

@@ 5.1 Datasets and Baselines
Table 1: Performance of different methods on different datasets. The solid lines separate different PEFT methods, while the dashed lines separate different ratios of trainable parameters. In each dashed-line area, the first row represents the original method, and the second row represents our method.
Due to the limitation of uploading test set results to the official website only twice a day, we use only the validation set results.
5.2 Evaluation Metrics
5.3 Implementation Details
We follow the same parameter setting for the classification task as BERT Devlin et al. (2019). The initialization weight method of the matrices $B$ and $A$ in LoRA and DoRA follows PiSSA Meng et al. (2024). For experiments, we did not use any hyper-parameter tuning, nor did we use MLNI trick (use the MLNI checkpoint instead of the pre-trained weights) to enhance the models’ performance. More details about the hyperparameters are available in Table 6 in Appendix C.
For different trainable parameter ratios, we select various layers of the LLM as trainable parameters to achieve different trainable ratios. For most PEFT methods, we select the bottom 1, 2, 3, and 4 layers to form four groups with different trainable ratios. In FISH-Tuning method, we select the bottom 5 layers as the parameter set and use the FISH Mask to choose the top-k parameters from this set, ensuring the same trainable ratios as the original PEFT method. For $(\text{IA})^{3}$, we select the bottom 3, 5, 8, and 10 layers to form four groups with different trainable ratios, while in the FISH-Tuning method, we use 12 layers because $(\text{IA})^{3}$ trains fewer parameters compared to other PEFT methods.
5.4 Experimental Results
5.4.1 Baselines
We conduct extensive experiments on the GLUE dataset and the BERT model to verify the effectiveness of FISH-Tuning. Table 1 presents the detailed performance of FISH-Tuning and baseline methods on the GLUE benchmark. Overall, the results demonstrate that integrating the FISH Mask into various PEFT methods consistently improves performance across multiple tasks, except for UniPELT.
In LoRA, FISH-Tuning consistently outperforms the original method. At a low parameter ratio of 0.0057%, while the improvement on MRPC and RTE is minimal, FISH-Tuning achieves higher scores on CoLA, SST-2, and STS-B, resulting in an overall average boost (from 68.45 to 68.90). As the trainable parameter ratio increases (e.g., to 0.0099% and 0.0184%), the improvements become even more significant, with average scores rising from 70.42 to 72.25 and from 72.11 to 72.91.
In DoRA, FISH-Tuning enhances performance on nearly all tasks, with average scores increasing by up to 2.75 points at a parameter ratio of 0.0142%. Notably, significant improvements on tasks like CoLA and RTE highlight the robustness of FISH-Tuning across different parameter scales.
In Adapter, at a 0.1389% trainable parameter ratio, the original Adapter method achieves an average score of 67.85, while FISH-Tuning raises the average to 73.25—a gain of over 5 points. This trend continues at higher parameter ratios (0.2760%, 0.4127%, and 0.5490%), confirming that FISH-Tuning not only enhances performance but also scales well with an increased trainable parameter ratio.
In Prefix-Tuning, integrating the FISH Mask generally leads to notable improvements at lower trainable parameter ratios. For example, at a 0.0439% parameter ratio, FISH-Tuning raises the CoLA score from 34.03 to 38.17, SST-2 from 83.83 to 88.07, and STS-B from 80.65 to 82.92—resulting in an overall average increase from 65.40 to 67.54. Similarly, at 0.0864% and 0.1289%, the average performance improves from 65.83 to 67.90 and from 66.13 to 67.93, respectively. However, at a higher parameter ratio of 0.1713%, although gains are observed on CoLA and SST-2 (e.g., CoLA increases from 35.80 to 40.94), notable declines on tasks such as RTE (from 65.34 down to 60.29) and WNLI (from 64.79 down to 59.15) lead to an overall average drop from 69.34 to 68.49. These findings suggest that while FISH integration in Prefix-Tuning is effective at lower parameter scales, its impact becomes less consistent as more parameters are tuned.
In $(\text{IA})^{3}$, FISH-Tuning yields mixed outcomes that appear to depend on the parameter ratio. At a low ratio of 0.0142%, FISH-Tuning consistently improves performance—boosting CoLA from 34.11 to 38.54, SST-2 from 87.16 to 89.11, and STS-B from 84.68 to 86.60, which increases the average score from 65.81 to 67.46. When the ratio is increased to 0.0227%, improvements in MRPC and STS-B are offset by slight drops in CoLA and RTE, resulting in a minimal average gain (67.91 to 67.99). At 0.0354%, decreases in CoLA and RTE cause the average to drop from 67.54 to 67.19. Notably, at 0.0439%, the integration of FISH Mask leads to a significant boost in CoLA (from 39.59 to 46.74) along with modest gains in RTE and SST-2, raising the average from 67.64 to 68.96. Overall, while FISH-Tuning can enhance $(\text{IA})^{3}$ performance under certain settings, its benefits are somewhat sensitive to the trainable parameter ratio.
In contrast, the integration of the FISH Mask in UniPELT does not yield consistent benefits. At a 0.0213% trainable parameter ratio, despite improvements in CoLA (increasing from 47.25 to 49.72) and RTE (from 60.65 to 64.26), there is a dramatic drop in WNLI—from 66.20 to 43.66—which pulls the average score down from 72.04 to 69.33. This adverse trend persists across higher parameter ratios: at 0.0411%, the average falls from 72.73 to 70.72; at 0.0610%, from 73.85 to 70.89; and at 0.0808%, from 72.72 to 70.92. We give our possible hypothesis for this experiment result in Limitations.
@@ 5.4.1 Baselines
Table 2: Performance of different tasks in GLUE. (M) means Modern BERT. (L) means LLaMA-3.2-1B.
5.4.2 Loss on Evaluation Datasets
We conduct an experiment on the loss scores using the evaluation dataset. The loss values are taken from the first solid-line area in Table 1. We visualize the loss value curves for the original PEFT method and FISH-Tuning method in Fig. 3. In this figure, the x-axis represents the training epochs, while the y-axis denotes the average evaluation loss scores across six datasets. From this figure, we observe that all points on the FISH-Tuning curve are lower than those on the original PEFT method curve. This result demonstrates that the FISH-Tuning method converges faster than the original PEFT method. It also explains why FISH-Tuning method achieves better results in Table 1.
Figure 3: Original LoRA method loss (Blue line) and the LoRA-FISH method loss (Orange line).
5.4.3 Results with other Pre-trained Models
Apart from the BERT model, we also compare FISH-Tuning with the original PEFT method using other trending pre-trained models, such as ModernBERT Warner et al. (2024) and LLaMA-3.2-1B Dubey et al. (2024). The prompt template for LLaMA-3.2-1B follows Table 1 in Zhong et al. (2023). Apart from the prompt setting, the rest of the settings are exactly the same as the BERT settings. The experiment results can be seen in Table 2.
In the ModernBERT experiments, we observe that integrating the FISH Mask into LoRA consistently improves performance across nearly all tasks. For instance, at a very low trainable parameter ratio of 0.0033%, FISH-Tuning boosts the CoLA score from 33.93 to 36.60 and enhances SST-2 from 83.37 to 86.24, leading to an overall average increase from 63.20 to 65.96. Similar trends are evident at slightly higher parameter ratios—at 0.0056% and 0.0080%, modest improvements in metrics such as RTE and MRPC are complemented by more pronounced gains in SST-2 and STS-B, yielding average scores rising from 63.50 to 65.75 and from 63.97 to 66.28, respectively. Even at 0.0103%, where the original LoRA attains an average of 66.17, FISH-Tuning further pushes the SST-2 and STS-B scores (up to 87.16 and 82.58, respectively), resulting in an average improvement to 66.81.
For LLaMA-3.2-1B, the benefits of FISH-Tuning are even more pronounced. At the lowest parameter ratio of 0.0007%, FISH-Tuning not only raises the CoLA score from 46.80 to 47.07 but also significantly improves MRPC and RTE (from 77.31 to 79.67 and 68.59 to 70.04, respectively), which lifts the overall average from 70.14 to 71.36. At a parameter ratio of 0.0010%, despite the strong baseline performance (average of 71.36), FISH-Tuning further enhances the performance to 72.76 by providing substantial gains on CoLA and RTE. With increasing parameter ratios of 0.0013% and 0.0017%, FISH-Tuning continues its steady performance improvements, achieving average scores of 72.96 and 73.03, respectively.
These results on ModernBERT and LLaMA-3.2-1B highlight the generalizability of FISH-Tuning across different pre-trained models, confirming that FISH-Tuning is effective and robust in enhancing the performance of PEFT methods beyond the standard BERT settings.
5.4.4 Contrastive Study
We further investigate the impact of the FISH Mask by conducting a contrastive study on the LoRA framework, as detailed in Table 5 in Appendix B. In this study, we compare the standard LoRA method with three variants: our proposed LoRA-FISH (the standard FISH-Tuning method), a variant where important parameters are selected at random (LoRA-FISH-rand), and another variant where the importance ordering is reversed (LoRA-FISH-rev). The parameter selection method of LoRA-FISH-rev is described as follows, replacing Equation 7:
At a low trainable parameter ratio of 0.0057%, LoRA-FISH improves the average score over the original LoRA—from 68.45 to 68.90. In contrast, while the random selection variant (LoRA-FISH-rand) attains a comparable average (68.74), the reverse-ordering strategy (LoRA-FISH-rev) leads to a noticeable drop, yielding an average score of only 66.71. As the parameter ratio increases to 0.0099%, the superiority of the FISH Mask selection becomes more pronounced. LoRA-FISH boosts the average score to 72.25 compared to 70.42 for the original method, whereas LoRA-FISH-rand and LoRA-FISH-rev obtain lower averages of 69.62 and 67.93, respectively.
A similar pattern is also observed at higher ratios. At 0.0142%, LoRA-FISH achieves an average score of 72.64, outperforming both the original LoRA (72.08) and the two contrastive variants (71.89 for the random selection and 71.86 for the reverse ordering). At the highest ratio tested (0.0184%), although both LoRA-FISH and its random counterpart yield comparable averages (72.91 and 72.92, respectively), the reverse ordering variant still falls slightly behind (72.65), and all three methods surpass the original method (72.11).
These findings reinforce that the specific parameter importance guided by the FISH Mask is significant: while randomly or reversed selection strategies may occasionally yield competitive performance at certain scales, they consistently fail to match the robust improvements delivered by FISH Mask approach.
6 Conclusion
In this paper, we propose FISH-Tuning, a new method that uses the selective PEFT method in the Addition-based and Reparameterization-based PEFT methods. More concretely, we integrate the FISH Mask into LoRA, DoRA, Adapter, Prefix-Tuning, and $(\text{IA})^{3}$. With the same ratio of trainable parameters, our method outperforms the original PEFT method most of the time. In future work, we will explore our approach in Computer Vision, Multi-modality, and Quantization.
Limitations
2) We use the LoRA method as an example. To achieve the same ratio of trainable parameters in FISH-Tuning, we can select different trainable layers and LoRA ranks. Tables 3 and 4 in Appendix A demonstrate our experiments on this. According to the results, we have not found any obvious pattern for determining how to select the suitable ranks and layers to achieve the best result.
Nevertheless, we are excited to see the huge potential of FISH-Tuning already demonstrated in existing experiments and look forward to more tests and suggestions from the community.
Ethics Statement
Acknowledgments
Appendix A LoRA Different Layer
@@
Table 3: Performance of different methods on different datasets. We set the LoRA rank to 1 and select different layers as trainable parameters.
@@
Table 4: Performance of different methods on different datasets. We set the trainable layers to 5 and select different LoRA ranks.
We select different trainable layers and LoRA ranks. The results in Tables 3 and 4.
Appendix B LoRA random and LoRA reverse selection
We compare the standard LoRA method with three variants: our proposed LoRA-FISH (the standard FISH-Tuning method), a variant where important parameters are selected at random(LoRA-FISH-rand), and another variant where the importance ordering is reversed (LoRA-FISH-rev). The results in Tables 5.
Method Trainable Parameters CoLA MRPC RTE SST-2 STS-B WNLI Avg
Original-LoRA 0.0057% 43.27 80.94 58.48 89.56 84.94 53.52 68.45
LoRA-FISH 0.0057% 44.28 80.25 58.48 90.48 86.41 53.52 68.90
LoRA-FISH-rand 0.0057% 46.18 78.38 58.12 89.56 86.69 53.52 68.74
LoRA-FISH-rev 0.0057% 38.91 74.80 58.12 88.99 85.91 53.52 66.71
[…]
LoRA-FISH 0.0099% 51.21 85.74 66.06 90.14 86.84 53.52 72.25
LoRA-FISH-rand 0.0099% 48.67 81.25 58.12 89.22 86.92 53.52 69.62
LoRA-FISH-rev 0.0099% 44.53 74.80 58.12 89.91 86.68 53.52 67.93
[…]
LoRA-FISH 0.0142% 53.58 85.56 65.70 90.71 86.79 53.52 72.64
LoRA-FISH-rand 0.0142% 52.07 82.24 66.43 90.14 86.95 53.52 71.89
LoRA-FISH-rev 0.0142% 48.68 84.03 66.43 91.17 87.33 53.52 71.86
[…]
LoRA-FISH 0.0184% 55.96 84.18 67.51 89.45 86.81 53.52 72.91
LoRA-FISH-rand 0.0184% 54.59 85.49 66.79 89.91 87.23 53.52 72.92
LoRA-FISH-rev 0.0184% 54.10 83.31 67.51 90.25 87.20 53.52 72.65
Table 5: Performance of different methods on different datasets. In each dashed-line area, the first row represents the original method, the second row represents our method, the third row represents the method where we randomly select the important parameters without using the FISH Mask, and the fourth row represents the method where we select the important parameters in reverse order compared to the second method.
Appendix C Hyperparameter settings of the experiment
@@
batch size 32
learning rate 5e-5, except WNLI 5e-6
epoch 400
optimizer Adam
early stop Yes
@@
warm up ratio 0.0
number of FISH Mask samples 128
dataset used for Fisher estimation train set
Prefix-Tuning’s prefix length $l$ 30
@@
LoRA dropout rate 0.0
Selected LoRA weights For BERT, we select $W_{Q}$, $W_{K}$, and $W_{V}$.
For ModernBERT and LLaMA-3.2-1B, we select $W_{O}$.
max sequence length details in Table 7

### v3

@@ Appendix B Tables of Precision Results
Table 6: Performance of different PEFT methods on GLUE dataset. The solid lines separate different PEFT methods, and the dashed lines separate different ratios of trainable parameters. In each dashed-line area, the first row represents the original method, and the second row represents our method.
@@
Table 7: Performance of different PEFT methods on GLUE dataset. (M) means ModernBERT-base. (L) means LLaMA-3.2-1B.
Method Trainable Parameters CoLA MRPC RTE SST-2 STS-B WNLI Avg
Original-LoRA 0.0057% 43.27 80.94 58.48 89.56 84.94 53.52 68.45
LoRA-FISH 0.0057% 44.28 80.25 58.48 90.48 86.41 53.52 68.90
LoRA-FISH-rand 0.0057% 46.18 78.38 58.12 89.56 86.69 53.52 68.74
LoRA-FISH-rev 0.0057% 38.91 74.80 58.12 88.99 85.91 53.52 66.71
[…]
LoRA-FISH 0.0099% 51.21 85.74 66.06 90.14 86.84 53.52 72.25
LoRA-FISH-rand 0.0099% 48.67 81.25 58.12 89.22 86.92 53.52 69.62
LoRA-FISH-rev 0.0099% 44.53 74.80 58.12 89.91 86.68 53.52 67.93
[…]
LoRA-FISH 0.0142% 53.58 85.56 65.70 90.71 86.79 53.52 72.64
LoRA-FISH-rand 0.0142% 52.07 82.24 66.43 90.14 86.95 53.52 71.89
LoRA-FISH-rev 0.0142% 48.68 84.03 66.43 91.17 87.33 53.52 71.86
[…]
LoRA-FISH 0.0184% 55.96 84.18 67.51 89.45 86.81 53.52 72.91
LoRA-FISH-rand 0.0184% 54.59 85.49 66.79 89.91 87.23 53.52 72.92
LoRA-FISH-rev 0.0184% 54.10 83.31 67.51 90.25 87.20 53.52 72.65
Table 8: Performance of different PEFT methods on GLUE dataset. In each dashed-line area, the first row represents the original method, the second row represents our method, the third row represents the method where we randomly select the important parameters without using the FISH Mask, and the fourth row represents the method where we select the important parameters in reverse order compared to the second method.
Method Ratio1 Ratio2 CoLA MRPC RTE SST-2 STS-B WNLI Avg
Original-LoRA 0.8116% 100% 60.85 88.26 70.40 91.86 88.77 53.52 75.61
[…]
LoRA-FISH-rand 0.4058% 50% 57.54 86.36 69.31 91.28 88.29 53.52 74.39
LoRA-FISH-rev 0.4058% 50% 58.56 80.35 61.37 91.97 87.68 53.52 72.24
[…]
LoRA-FISH-rand 0.2029% 25% 56.29 85.45 71.12 91.17 87.99 53.52 74.26
LoRA-FISH-rev 0.2029% 25% 51.62 78.27 58.84 91.40 87.00 53.52 70.11
[…]
LoRA-FISH-rand 0.0812% 10% 52.89 81.64 63.18 91.17 87.91 53.52 71.72
LoRA-FISH-rev 0.0812% 10% 49.95 74.80 58.12 90.14 86.27 53.52 68.80
[…]
LoRA-FISH-rand 0.0406% 5% 49.11 80.90 58.84 91.28 87.84 53.52 70.25
LoRA-FISH-rev 0.0406% 5% 39.18 74.80 58.12 88.42 85.00 53.52 66.51
[…]
LoRA-FISH-rand 0.0081% 1% 34.44 74.80 58.48 89.68 85.46 53.52 66.07
LoRA-FISH-rev 0.0081% 1% 25.12 74.80 58.12 83.83 76.06 53.52 61.91
[…]
LoRA-FISH-rand 0.0041% 0.5% 34.55 74.80 58.12 87.61 82.19 53.52 65.13
LoRA-FISH-rev 0.0041% 0.5% 18.81 74.80 58.12 81.77 71.12 53.52 59.69
Table 9: Performance of different PEFT methods on GLUE dataset. The "Ratio1" column represents the ratio of trainable parameters to the total number of parameters, and the "Ratio2" column represents the ratio of trainable parameters to the total number of parameters in LoRA component.
Appendix C Different Layers and Ranks for LoRA
@@
Table 10: Performance of different PEFT methods on GLUE dataset. We set the LoRA rank to 1 and select different layers as trainable parameters.
@@
Table 11: Performance of different PEFT methods on GLUE dataset. We set the trainable layers to 5 and select different LoRA ranks.
We select different trainable layers and LoRA ranks. The results in Tables 10 and 11.
Appendix D Hyperparameter Settings of the Experiment
@@
batch size 32 in classification task
4 in generation task
gradient accumu- 1 in classification task
lation steps 4 in generation task
learning rate 5e-5 in classification task
5e-6 in WNLI
2e-5 in generation task
epoch 400 in classification task
10 in generation task
early stop Yes in classification task
No in generation task
optimizer AdamW
@@
warm up ratio 0.0 in classification task
0.03 in generation task
lr scheduler linear in classification task
type cosine in generation task
number of FISH 128 in classification task
Mask samples 256 in generation task
dataset used for train set
Fisher estimation
Prefix-Tuning’s 30
prefix length $l$
LoRA rank details in Table 13
@@
LoRA dropout 0.1 in Qwen
rate 0.0 in other model
Selected LoRA details in Table 14
weights
distributed type zero-3 offload
(generation task)
Nvidia GPU 4 * A100 80G for Qwen
8 * A5000 24G for others
max sequence details in Table 15
length
@@
Task-fixing the proportion LoRA Rank Value
of trainable parameters
BERT-base-cased 1
ModernBERT-base 1
LLaMA-3.2-1B 1
Task-fixing the number LoRA Rank Value
of trainable layers
BERT 16
Phi-4-mini-instruct 64
Qwen2.5-7B 64
Table 13: The different LoRA Rank value in different tasks.
Pre-trained model Selected LoRA Weights
BERT-base-cased $W_{Q}$, $W_{K}$, $W_{V}$
ModernBERT-base $W_{O}$
LLaMA-3.2-1B $W_{O}$
Phi-4-mini-instruct $W_{Q}$, $W_{K}$, $W_{V}$
Qwen2.5-7B $W_{Q}$, $W_{K}$, $W_{V}$
Table 14: The different selected LoRA weights in different pre-trained models.
@@
Appendix E Evaluation Metrics
GLUE is a multi-task benchmark that contains 10 datasets for LLM evaluation. In our experiment, we select only the CoLA Warstadt et al. (2018), MRPC Dolan and Brockett (2005), RTE, SST-2 Socher et al. (2013), STS-B Cer et al. (2017), and WNLI Levesque et al. (2012) datasets because the remaining datasets contain too much text and require excessive training time. Then we calculate the average score of these six datasets. Due to the limitation of uploading test set results to the official website only twice a day, we use only the validation set results.
GSM8K is a dataset consisting of 8.5K high-quality grade school math problems designed to support multi-step reasoning tasks. It includes detailed step-by-step solutions for each problem, aiming to enhance the mathematical reasoning capabilities of machine learning models.
Appendix F Pictures of Specific Results for Generation Task (Qwen2.5-7B)
Fig. 8 and Fig. 9 show the results of different epochs and ratios in Qwen2.5-7B model.
Figure 8: Performance of FISH-Tuning method in different proportion of Trainable Parameters on the GSM8K dataset using Qwen2.5-7B model for different training epochs.
Figure 9: LoRA-FISH converges faster compared to LoRA-FISH-random and LoRA-FISH-reverse using Qwen2.5-7B model for different proportion of Trainable Parameters.
Appendix G Pictures of Specific Results for Generation Task (Phi-4-mini-instruct)
Fig. 10 and Fig. 11 show the results of different epochs and ratios in Phi-4-mini-instruct model.
Figure 10: Performance of FISH-Tuning method in different proportion of Trainable Parameters on the GSM8K dataset using Phi-4-mini-instruct model for different training epochs.
Figure 11: LoRA-FISH converges faster compared to LoRA-FISH-random and LoRA-FISH-reverse using Phi-4-mini-instruct model for different proportion of Trainable Parameters.
Figure 12: Original FISH Mask method (left) and original LoRA method (mid) and our FISH-Tuning method (right).